# One-Shot Real-World Demonstration Synthesis for Scalable Bimanual Manipulation

### v1

@@ Abstract
Learning dexterous bimanual manipulation policies critically depends on large-scale, high-quality demonstrations, yet current paradigms face inherent trade-offs: teleoperation provides physically grounded data but is prohibitively labor-intensive, while simulation-based synthesis scales efficiently but suffers from sim-to-real gaps. We present BiDemoSyn, a framework that synthesizes contact-rich, physically feasible bimanual demonstrations from a single real-world example. The key idea is to decompose tasks into invariant coordination blocks and variable, object-dependent adjustments, then adapt them through vision-guided alignment and lightweight trajectory optimization. This enables the generation of thousands of diverse and feasible demonstrations within several hour, without repeated teleoperation or reliance on imperfect simulation. Across six dual-arm tasks, we show that policies trained on BiDemoSyn data generalize robustly to novel object poses and shapes, significantly outperforming recent baselines. By bridging the gap between efficiency and real-world fidelity, BiDemoSyn provides a scalable path toward practical imitation learning for complex bimanual manipulation without compromising physical grounding.
1 Introduction
@@
Recent advances in robot manipulation have been propelled by data-driven imitation learning, where visuomotor policies trained on large-scale demonstration datasets enable dexterous single-arm and bimanual tasks previously deemed intractable. Methods like Action Chunking Transformer (ACT) zhao2023learning, Diffusion Policy (DP) chi2023diffusion and $\pi_{0}$ black2024pi0 exemplify this trend, achieving remarkable performance in contact-rich scenarios such as continuously pushing, battery assembly, and coordinated garment folding. These successes, however, hinge on a critical yet underappreciated premise: the availability of diverse, high-quality demonstrations, which are typically collected via expert teleoperation. As the community shifts toward more complex long-horizon bimanual tasks, the data scalability bottleneck becomes increasingly apparent. Even those largest manipulation datasets like RH20T fang2024rh20t, DROID khazatsky2024droid, Open X-Embodiment o2024open and AgiBot World bu2025agibot struggle to provide sufficient geometric and kinematic diversity for robust generalization, underscoring a pivotal question: how to scale real-world demonstration collection without compromising physical fidelity?
[…]
@@
The answer, unfortunately, lies in a fundamental trade-off. Existing data acquisition methods fall into two paradigms, each with critical shortcomings. On one hand, human-operated systems (e.g., the ALOHA series zhao2023learning; fu2024mobile; zhao2024aloha) yield physically-accurate trajectories but demand prohibitive expertise, rendering them impractical for scaling to long-horizon tasks. On the other hand, simulation-based methods (e.g., MimicGen mandlekar2023mimicgen, RoboGen wang2024robogen, RoboCasa nasiriany2024robocasa, and DexMimicGen jiang2024dexmimicgen) bypass human labor through programmatic data generation utilizing massive 3D assets, yet discrepancies in contact dynamics and visual rendering inevitably plague real-world deployment bharadhwaj2024position; mccarthy2024towards. Thus, the tension between data quality (grounded in reality) and quantity (enabled by automation) persists as a key barrier to scalable imitation learning, particularly for contact-rich bimanual tasks requiring precise dual-arm coordination.
@@
To bridge this gap, we present BiDemoSyn that synthesizes real-world bimanual demonstrations from a single exemplar, by algorithmically amplifying task semantics through a hierarchical decomposition of invariance and adaptability. Our key insight lies in recognizing that complex manipulation tasks exhibit a duality: while certain motion primitives (e.g., stable grasp sequences, dual-arm synchronization) remain invariant across instances, others (e.g., trajectory segments adapting to object shifting) require dynamic adjustments to contextual variations. Through a three-stage process, BiDemoSyn first deconstructs the initial demonstration into invariant primitives and geometry-sensitive components, then leverages visual perception to spatially align these elements with novel scenes, enabling generalization across object geometries and workspace layouts with minor human effort (primarily involving reinitializing manipulated objects). Finally, a physics-aware optimizer jointly modulates dual-arm trajectories, injecting diversity into adaptable components while enforcing real-world constraints on collision avoidance and kinematic coordination. This integrated approach achieves one-to-many demonstration synthesis: a single exemplar spawns hundreds of trajectories that preserve task intent while adapting to unseen configurations, all grounded in real-world dynamics rather than simulated proxies. Examples are shown in Fig. 1 and Fig. 2.
@@
Unlike prior simulation-based methods mandlekar2023mimicgen; wang2024robogen; nasiriany2024robocasa; jiang2024dexmimicgen, BiDemoSyn operates entirely within the physical domain, ensuring synthesized data inherits the fidelity of human demonstrations. By unifying algorithmic scalability with physical realism, our framework empowers policies to transcend the limitations of narrow manually-collected data distributions. Experiments on real dual-arm platforms validate the effectiveness of BiDemoSyn. And visuomotor policies trained with our synthesized demonstrations can achieve superior success rates and generalization in tasks requiring precise contact coordination (e.g., small hole insertion, rotation angle control, purposeful handover and articulated object manipulation).
@@ 2 Related Works
Bimanual Robotic Manipulation. Prior literatures primarily target task-specific challenges (such as cloth folding colome2018dimensionality; weng2022fabricflownet; canberk2023cloth, untangling grannen2021untangling; peng2024tiebot, untwisting lin2024twisting; bahety2024screwmimic, and handover huang2023dynamic; li2023efficient) through handcrafted controllers or narrow skill libraries, limiting generalization due to rigid motion primitives and excessive assumptions. General methods xie2020deep; chitnis2020efficient; chen2022towards; chen2023bi; hartmann2022long; zhaodual2023afford either rely on predefined hierarchical arm roles (e.g., leader-follower krebs2022bimanual; liu2022robot; grotz2024peract2 and stabilizer-actor grannen2023stabilize; liu2024voxact) or hinge generalization entirely on dataset diversity, both of which are ultimately bottlenecked by the impractical costs of data acquisition. Recently, ALOHA series zhao2023learning; fu2024mobile; zhao2024aloha have revolutionized bimanual manipulation by dexterous low-cost teleoperation. Meanwhile, the visuomotor policy learning paradigm chi2023diffusion; ze2024dp3; yang2024equibot based on diffusion models ho2020denoising; song2021denoising has largely expanded the action modeling space and data utilization efficiency. Their followers team2024octo; kim2024openvla; liu2025rdt; black2024pi0; pertsch2025fast; lin2025data expect to train universal policies using large-scale teleoperation data but face scalability barriers from human effort. To improve reachability and dexterity, some studies use specialized hardware (e.g., multi-finger hands wang2024dexcap; shaw2024bimanual; fu2024humanplus; cheng2024open or tactile sensors lin2024learning; chen2024arcap), which complicate real-world adoption. In contrast, we utilize a fixed-base dual-arm platforrm with parallel grippers, synthesizing diverse demonstrations from a single example. By optimizing coordination as dynamic constraints rather than predefined hierarchies, we bypass hardware specificity and data scalability limitations, enabling task-agnostic policies grounded in physical feasibility.
@@
Bimanual Demonstration Collection. The dominant path for acquiring bimanual demonstrations is human teleoperation khazatsky2024droid; bu2025agibot, which delivers high-quality data but suffers from prohibitive scalability costs. To alleviate it, two alternative routes have emerged: simulation-based synthesis garrett2024skillmimicgen; hua2024gensim2; liang2024make; yang2024physcene; wang2024gensim and learning from human videos ponimatkin20256d; ye2025video2policy; zhao2025taste; kareer2024egomimic; bharadhwaj2024track2act. The former, exemplified by MimicGen mandlekar2023mimicgen and DexMimicGen jiang2024dexmimicgen, augments a few demonstrations by generating variations in simulation using geometric transformations and kinematic constraints. Analogously, RoboGen wang2024robogen and RoboCasa nasiriany2024robocasa leverage 3D assets to procedurally generate full-scene manipulation data, yet such methods inevitably inherit sim-to-real gaps, from unrealistic textures to unphysical dynamics. The latter extracts bimanual hand-object manipulation from egocentric videos zhan2024oakink2; liu2024taco; grauman2024ego and maps them to dual arms via motion retargeting li2024okami; kerr2024robot; chen2024object; chen2024vividex or non-privileged representations (e.g., keypoints papagiannis2024rx; gao2024bi; wen2023any, affordances ju2024robo; nasiriany2024rt and correspondences peng2024tiebot; ko2024learning; zhang2024one). While human videos are abundant and low-cost, large morphological mismatches between humanity and robotic embodiments often need heuristic translation rules and hinder direct applicability. Our work navigates these trade-offs by synthesizing demonstrations directly in real-world. Given a single exemplar, we diversify trajectories through vision-based adaptation and physics-compliant optimization. This balances the fidelity and scalability, while avoiding retargeting hurdles, thereby enabling practical and scalable data acquisition for contact-rich bimanual tasks.
@@
Several concurrent works exhibit notable limitations. DemoGen xue2025demogen and YOTO zhou2025you diversify demonstrations via 3D editing of the initial single-view point cloud, but perspective ambiguities induce visual artifacts in synthesized data. ODIL wang2025one avoids editing by segmenting objects for visual servoing and iteratively replanning trajectories, which is cumbersome for scalability. MoMaGen li2025momagen is a long-horizon mobile+bimanual demonstrations generator yet mainly focusing on the simulation domain. Our BiDemoSyn sidesteps these issues: synthesizing trajectories directly in real-world ensures visual-physical consistency without manual edits or robot replaying.
3 Preliminaries
@@
Problem Formulation. Bimanual imitation learning aims to train visuomotor policies $\pi_{\theta}(\textbf{o}_{t})\rightarrow\textbf{a}_{t}$, where $\textbf{o}_{t}$ denotes multimodal observations (e.g., RGB images, joint states or point clouds) and $\textbf{a}_{t}$ represents dual-arm actions. While architectures like ACT zhao2023learning and DP chi2023diffusion enhance policy expressivity, their generalization critically depends on expert demonstrations $\mathcal{D}\!=\!\{\tau_{i}\}_{i=0}^{N}$ that densely span the feasible state-action manifold. Existing approaches either collect $\mathcal{D}$ via labor-intensive teleoperation or generate data in simulation, both failing to balance real-world fidelity and scalability. We address this gap by proposing the following problem:
@@
Visuomotor Policy Learning. Our $\pi_{\theta}$ maps observations $\textbf{o}_{t}$ to dual-arm actions $\textbf{a}_{t}\!=\!\left[\textbf{a}_{t}^{L},\textbf{a}_{t}^{R}\right]\!\in\!\mathbb{R}^{2d}$, where $\textbf{a}_{t}^{L}$, $\textbf{a}_{t}^{R}$ denotes left/right end-effector poses. We adapt a diffusion policy model chi2023diffusion; ze2024dp3; yang2024equibot to generate action sequences $\textbf{A}\!=\!\left[\textbf{a}_{t:{t+H}}^{L},\textbf{a}_{t:{t+H}}^{R}\right]\!\in\!\mathbb{R}^{2H\times d}$ over a horizon $H$. The diffusion process iteratively denoises a noisy action sequence $\textbf{A}_{k}$ toward kinematic feasibility over $K$ steps:
@@
where $\alpha_{k}$ follows a cosine schedule nichol2021improved. And $\mathbf{\Sigma}\!\in\!\mathbb{R}^{2H\times 2H}$ is a block-diagonal covariance matrix $\left[\begin{array}[]{cc}\mathbf{\Sigma}_{L}&\rho\mathbf{\Sigma}_{LR}\\ \rho\mathbf{\Sigma}^{\top}_{LR}&\mathbf{\Sigma}_{R}\end{array}\right]$ encoding arm coordination priors, where $\mathbf{\Sigma}_{L}$, $\mathbf{\Sigma}_{R}$ governing temporal smoothness per arm, $\mathbf{\Sigma}_{LR}$ modeling inter-arm dependencies, and $\rho\!\in\![0,1]$ controlling coordination strength. Then, a denoiser $\mu_{\theta}$ processes the noised sequence conditioned on visual observations $\textbf{o}_{t}$ via:
@@
where $\mathsf{E}(\textbf{o}_{t})$ is a vision encoder (ResNet he2016deep or PointNet qi2017pointnet) extracting latent features. The $\texttt{UNet}_{\theta}$ employs asymmetric layers with early stages processing each arm independently, while deeper layers fusing bilateral coordinated features. Training minimizes:
@@
Figure 2: The overview of BiDemoSyn. It consists of three stages (e.g., deconstruction, alignment, and optimization) based on a given demonstration. Then, we can apply our method to complete data collection efficiently and conveniently in real-world. It is best to zoom in to view the details.
4 Methodology
Here, we detail BiDemoSyn for synthesizing bimanual demonstrations $\mathcal{D}_{syn}$ from a single exemplar $\tau$ (teleoperated or kinesthetically guided) with three stages: Sec. 4.1 Deconstruction of One-Shot Teaching, which extracts invariant patterns and adaptable primitives; Sec. 4.2 Vision-Based Initial Frame Alignment, enabling generalization across geometric variations via efficient scene perception; and Sec. 4.3 Trajectory Modulation and Optimization, ensuring physical feasibility through hierarchical kinematic constraints. Each stage sequentially addresses scalability, adaptability, and real-world fidelity to bridge the gap between data efficiency and policy robustness. The overall framework is shown and illustrated in Fig. 2.
4.1 Deconstruction of One-Shot Teaching
Given a single demonstration $\tau\!=\!\{\textbf{o}_{t},\textbf{a}_{t}\}_{t=0}^{T}$, we decompose it into a bimanual execution blocks set $\{\mathcal{B}_{i}\}_{i=1}^{n}$, where each block $\mathcal{B}_{i}\!=\!(\textbf{s}_{i}^{L},\textbf{s}_{i}^{R})$ represents a discrete phase of dual-arm interaction. Here, $\textbf{s}_{i}^{L},\textbf{s}_{i}^{R}\!\in\!\mathbb{R}^{d}$ denote the left/right arm state sequences (including end-effector 6-DoF poses and gripper status) within block $\mathcal{B}_{i}$. Blocks are categorized based on motion coordination. For the single-arm motion, one arm exhibits significant state transitions while the other remains static. Formally, for threshold $\delta$ (minimal motion saliency) and $\zeta$ (static tolerance):
$\centering(\left\|\textbf{s}_{i,e}^{L}-\textbf{s}_{i,s}^{L}\right\|\!\geq\!\delta\wedge\left\|\textbf{s}_{i,e}^{R}-\textbf{s}_{i,s}^{R}\right\|\!\leq\!\zeta)\;\vee\;(\left\|\textbf{s}_{i,e}^{R}-\textbf{s}_{i,s}^{R}\right\|\!\geq\!\delta\wedge\left\|\textbf{s}_{i,e}^{L}-\textbf{s}_{i,s}^{L}\right\|\!\leq\!\zeta).\@add@centering$ (4)
While, for the dual-arm coordination, both arms synchronously or asynchronously adjust states:

### v2

@@ Abstract
Learning dexterous bimanual manipulation policies critically depends on large-scale, high-quality demonstrations, yet current paradigms face inherent trade-offs: teleoperation provides physically grounded data but is prohibitively labor-intensive, while simulation-based synthesis scales efficiently but suffers from sim-to-real gaps. We present BiDemoSyn, a framework that synthesizes contact-rich, physically feasible bimanual demonstrations from a single real-world example. The key idea is to decompose tasks into invariant coordination blocks and variable, object-dependent adjustments, then adapt them through vision-guided alignment and lightweight trajectory optimization. This enables the generation of thousands of diverse and feasible demonstrations within several hour, without repeated teleoperation or reliance on imperfect simulation. Across six dual-arm tasks, we show that policies trained on BiDemoSyn data generalize robustly to novel object poses and shapes, significantly outperforming recent strong baselines. Beyond the one-shot setting, BiDemoSyn naturally extends to few-shot-based synthesis, improving object-level diversity and out-of-distribution generalization while maintaining strong data efficiency. Moreover, policies trained on BiDemoSyn data exhibit zero-shot cross-embodiment transfer to new robotic platforms, enabled by object-centric observations and a simplified 6-DoF end-effector action representation that decouples policies from embodiment-specific dynamics. By bridging the gap between efficiency and real-world fidelity, BiDemoSyn provides a scalable path toward practical imitation learning for complex bimanual manipulation without compromising physical grounding.
I Introduction
@@
Recent advances in robot manipulation have been propelled by data-driven imitation learning, where visuomotor policies trained on large-scale demonstration datasets enable dexterous single-arm and bimanual tasks previously deemed intractable. Methods like ACT [103], DP [16] and $\pi_{0}$ [5] exemplify this trend, achieving remarkable performance in contact-rich scenarios such as continuously pushing, battery assembly, and coordinated garment folding. These successes, however, hinge on a critical yet underappreciated premise: the availability of diverse, high-quality demonstrations, which are typically collected via expert teleoperation. As the community shifts toward more complex long-horizon bimanual tasks, the data scalability bottleneck becomes increasingly apparent. Even those largest manipulation datasets like RH20T [22], DROID [44], Open X-Embodiment [68] and AgiBot World [6] struggle to provide sufficient geometric and kinematic diversity for robust generalization, underscoring a pivotal question: how to scale real-world bimanual demonstration collection without compromising physical fidelity?
@@
The answer, unfortunately, lies in a fundamental trade-off. Existing data acquisition methods fall into two paradigms, each with critical shortcomings. On one hand, human-operated systems (e.g., the ALOHA series [103, 26, 104]) yield physically-accurate trajectories but demand prohibitive expertise, rendering them impractical for scaling to long-horizon tasks. On the other hand, simulation-based methods (e.g., MimicGen [62, 40], RoboGen [85], and RoboCasa [66]) bypass human labor through programmatic data generation utilizing massive 3D assets, yet discrepancies in contact dynamics and visual rendering inevitably plague real-world deployment [3, 64]. Thus, the tension between data quality (grounded in reality) and quantity (enabled by automation) persists as a key barrier to scalable imitation learning, particularly for contact-rich bimanual tasks requiring precise dual-arm coordination.
@@
To bridge this gap, we present BiDemoSyn that synthesizes real-world bimanual demonstrations from a single exemplar, by algorithmically amplifying task semantics through a hierarchical decomposition of invariance and adaptability (refer Fig. LABEL:BiDemoSyn and Fig. 1). Our motivation is inherited from the well-established single-arm one-shot imitation learning [63, 60, 4, 100, 21, 106]. The key insight lies in recognizing that complex manipulation tasks exhibit a duality: while certain motion primitives (e.g., stable grasp sequences, dual-arm synchronization) remain invariant across instances, others (e.g., trajectory segments adapting to object shifts) require dynamic adjustments to contextual variations. Specifically, through a three-stage process, BiDemoSyn first deconstructs the initial demonstration into invariant primitives and geometry-sensitive components. Then, it leverages visual perception to spatially align these elements in novel scenes, enabling generalization across object geometries and workspace layouts with minor effort (primarily involving reinitializing objects). Finally, a physics-aware optimizer jointly modulates dual-arm trajectories, injecting diversity into adaptable components while enforcing real-world constraints on collision avoidance and kinematic coordination. This integrated approach achieves One-to-Many demonstration synthesis: a single exemplar spawns hundreds of trajectories that preserve task intent while adapting to unseen configurations, all grounded in real-world dynamics rather than simulated proxies.
@@
Unlike prior simulation-based methods [62, 85, 66, 40], BiDemoSyn operates entirely in the physical domain, ensuring synthesized data inherits the fidelity of human demonstrations. By unifying algorithmic scalability with physical realism, our framework empowers policies to transcend the limitations of narrow manually-collected data distributions. Experiments on real dual-arm platforms validate the effectiveness of BiDemoSyn. Visuomotor policies trained with synthesized demonstrations can achieve superior success rates and generalization in tasks requiring precise contact coordination (e.g., small hole insertion, rotation angle control, purposeful handover and articulated object manipulation).
@@ II Related Works
Bimanual Robotic Manipulation. Prior literatures primarily target task-specific challenges (such as cloth folding [20, 88, 7], untangling [29, 70], untwisting [54, 1], and handover [37, 51]) through handcrafted controllers or narrow skill libraries, limiting generalization due to rigid motion primitives and excessive assumptions. General methods [90, 18, 12, 10, 33, 105] either rely on predefined hierarchical arm roles (e.g., leader-follower [48, 57, 32] and stabilizer-actor [30, 56]) or hinge generalization entirely on dataset diversity, both of which are ultimately bottlenecked by the impractical costs of data acquisition. Recently, ALOHA series [103, 26, 104] have revolutionized bimanual manipulation by dexterous low-cost teleoperation. Meanwhile, the visuomotor policy learning paradigm [16, 98, 94] based on diffusion models [35, 79] has largely expanded the action modeling space and data utilization efficiency. Their followers [80, 45, 58, 5, 71, 53] expect to train universal policies using large-scale teleoperation data but face scalability barriers from human effort. To improve reachability and dexterity, some studies use specialized hardware (e.g., multi-finger hands [82, 76, 25, 15] or tactile sensors [55, 9]), which complicate real-world adoption. In contrast, we utilize a fixed-base dual-arm platforrm with parallel grippers, synthesizing diverse demonstrations from a single example. By optimizing coordination as dynamic constraints rather than predefined hierarchies, we bypass hardware specificity and data scalability limitations, enabling task-agnostic policies grounded in physical feasibility.
@@
Bimanual Demonstration Collection. The dominant path for acquiring bimanual demonstrations is human teleoperation [44, 6], which delivers high-quality data but suffers from prohibitive scalability costs. To alleviate it, two alternative routes have emerged: simulation-based synthesis [28, 36, 52, 96, 83] and learning from human videos [72, 97, 102, 42, 2]. The former, exemplified by MimicGen [62] and DexMimicGen [40], augments a few demonstrations by generating variations in simulation using geometric transformations and kinematic constraints. Analogously, RoboGen [85] and RoboCasa [66] leverage 3D assets to procedurally generate full-scene manipulation data, yet such methods inevitably inherit sim-to-real gaps, from unrealistic textures to unphysical dynamics. The latter extracts bimanual hand-object manipulation from egocentric videos [99, 59, 31] and maps them to dual arms via motion retargeting [50, 43, 11, 13] or non-privileged representations (e.g., keypoints [69, 27, 87], affordances [41, 65] and correspondences [70, 46, 101]). While human videos are abundant and low-cost, large morphological mismatches between humanity and robotic embodiments often need heuristic translation rules and hinder direct applicability. Our work navigates these trade-offs by synthesizing demonstrations directly in real-world. Given a single exemplar, we diversify trajectories through vision-based adaptation and physics-compliant optimization. This balances the fidelity and scalability, while avoiding retargeting hurdles, thereby enabling practical and scalable data acquisition for contact-rich bimanual tasks.
@@
Several concurrent works exhibit notable limitations. DemoGen [93] and YOTO [107] diversify demonstrations via 3D editing of the initial single-view point cloud, but perspective ambiguities induce visual artifacts in synthesized data. ODIL [84] avoids editing by segmenting objects for visual servoing and iteratively replanning trajectories, which is cumbersome for scalability. MoMaGen [49] is a long-horizon mobile bimanual demonstrations generator yet mainly focusing on the simulation. Our BiDemoSyn sidesteps these issues: synthesizing trajectories directly in real-world ensures visual-physical consistency without manual edits or robot replaying.
III Preliminaries
@@
Problem Formulation. Bimanual imitation learning aims to train visuomotor policies $\pi_{\theta}(\textbf{o}_{t})\rightarrow\textbf{a}_{t}$, where $\textbf{o}_{t}$ denotes multimodal observations (e.g., RGB images, joint states or point clouds) and $\textbf{a}_{t}$ represents dual-arm actions. While architectures like ACT [103] and DP [16] enhance policy expressivity, their generalization critically depends on expert demonstrations $\mathcal{D}\!=\!\{\tau_{i}\}_{i=0}^{N}$ that densely span the feasible state-action manifold. Existing approaches either collect $\mathcal{D}$ via labor-intensive teleoperation or generate data in simulation, both failing to balance real-world fidelity and scalability. We address this gap by proposing the following problem:
@@
Visuomotor Policy Learning. Our $\pi_{\theta}$ maps observations $\textbf{o}_{t}$ to dual-arm actions $\textbf{a}_{t}\!=\!\left[\textbf{a}_{t}^{L},\textbf{a}_{t}^{R}\right]\!\in\!\mathbb{R}^{2d}$, where $\textbf{a}_{t}^{L}$, $\textbf{a}_{t}^{R}$ denotes left/right end-effector poses. We adapt a diffusion policy model [16, 98, 94] to generate action sequences $\textbf{A}\!=\!\left[\textbf{a}_{t:{t+H}}^{L},\textbf{a}_{t:{t+H}}^{R}\right]\!\in\!\mathbb{R}^{2H\times d}$ over a horizon $H$. The diffusion process iteratively denoises a noisy action sequence $\textbf{A}_{k}$ toward kinematic feasibility over $K$ steps:
@@
where $\alpha_{k}$ follows a cosine schedule [67]. And $\mathbf{\Sigma}\!\in\!\mathbb{R}^{2H\times 2H}$ is a block-diagonal covariance matrix $\left[\begin{array}[]{cc}\mathbf{\Sigma}_{L}&\rho\mathbf{\Sigma}_{LR}\\ \rho\mathbf{\Sigma}^{\top}_{LR}&\mathbf{\Sigma}_{R}\end{array}\right]$ encoding arm coordination priors, where $\mathbf{\Sigma}_{L}$, $\mathbf{\Sigma}_{R}$ governing temporal smoothness per arm, $\mathbf{\Sigma}_{LR}$ modeling inter-arm dependencies, and $\rho\!\in\![0,1]$ controlling coordination strength. Then, a denoiser $\mu_{\theta}$ processes the noised sequence conditioned on visual observations $\textbf{o}_{t}$ via:
@@
where $\mathsf{E}(\textbf{o}_{t})$ is a vision encoder (ResNet [34] or PointNet [73]) extracting latent features. The $\texttt{UNet}_{\theta}$ employs asymmetric layers with early stages processing each arm independently, while deeper layers fusing bilateral coordinated features. Training minimizes:
@@
Figure 1: The overview of BiDemoSyn. It consists of three stages (e.g., deconstruction, alignment, and optimization) based on a given demonstration. Then, we can apply our method to complete data collection efficiently and conveniently in real-world.
IV Methodology
Here, we detail BiDemoSyn for synthesizing bimanual demonstrations $\mathcal{D}_{syn}$ from a single exemplar $\tau$ with three compact stages (see Fig. 1): Sec. IV-A Deconstruction of One-Shot Teaching, which extracts invariant patterns and adaptable primitives; Sec. IV-B Vision-based Initial Frame Alignment, enabling generalization across geometric variations via efficient scene perception; and Sec. IV-C Trajectory Modulation and Optimization, ensuring physical feasibility via hierarchical kinematic constraints. Each stage sequentially addresses scalability, adaptability, and real-world fidelity to bridge the gap between data efficiency and policy robustness.
IV-A Deconstruction of One-Shot Teaching
Given a demonstration $\tau\!=\!\{\textbf{o}_{t},\textbf{a}_{t}\}_{t=0}^{T}$, we decompose it into a bimanual execution blocks set $\{\mathcal{B}_{i}\}_{i=1}^{n}$, where each block $\mathcal{B}_{i}\!=\!(\textbf{s}_{i}^{L},\textbf{s}_{i}^{R})$ represents a discrete phase of dual-arm interaction. Here, $\textbf{s}_{i}^{L},\textbf{s}_{i}^{R}\!\in\!\mathbb{R}^{d}$ denote the left/right arm state sequences (e.g., end-effector 6-DoF poses and gripper status) within block $\mathcal{B}_{i}$. Blocks are categorized based on motion coordination. For single-arm motion, one arm exhibits significant transitions while the other remains static. Formally, for threshold $\delta$ (minimal motion saliency) and $\zeta$ (static tolerance):
$(\left\|\textbf{s}_{i,e}^{L}-\textbf{s}_{i,s}^{L}\right\|\!\geq\!\delta\wedge\left\|\textbf{s}_{i,e}^{R}-\textbf{s}_{i,s}^{R}\right\|\!\leq\!\zeta)\;\vee\;$ (4)
$(\left\|\textbf{s}_{i,e}^{R}-\textbf{s}_{i,s}^{R}\right\|\!\geq\!\delta\wedge\left\|\textbf{s}_{i,e}^{L}-\textbf{s}_{i,s}^{L}\right\|\!\leq\!\zeta).$
While, for dual-arm coordination, both arms synchronously or asynchronously adjust states: